### v1

# Towards Robust Real-World Multivariate Time Series Forecasting: A Unified Framework for Dependency, Asynchrony, and Missingness

Abstract
Real-world time series data are inherently multivariate, often exhibiting complex inter-channel dependencies. Each channel is typically sampled at its own period and is prone to missing values due to various practical and operational constraints. These characteristics pose fundamental challenges related to channel dependency, sampling asynchrony, and missingness, all of which must be addressed to enable robust and reliable forecasting in practical settings. However, most existing architectures are built on oversimplified assumptions, such as identical sampling periods across channels and fully observed inputs at test time, which often do not hold in real-world scenarios. To bridge this gap, we propose ChannelTokenFormer, a Transformer-based forecasting model with a flexible architecture designed to explicitly capture cross-channel interactions, accommodate channel-wise asynchronous sampling, and effectively handle missing values. Extensive experiments on three benchmark datasets modified to reflect practical settings, along with one real-world industrial dataset, demonstrate the superior robustness and accuracy of ChannelTokenFormer under challenging real-world conditions.
1 Introduction
Accurate time series forecasting is critical in real-world applications such as industrial monitoring [17, 54], energy systems [46, 32, 14], and healthcare [12, 37], where predictive insights directly impact operational safety, resource efficiency, and long-term outcomes. This task is conventionally formulated under assumptions of fully observed data and equal sampling periods across channels. In contrast to these assumptions, time series data in some real-world applications are often far more complex than what existing models assume, as they typically consist of multivariate signals originating from heterogeneous sources such as sensors measuring temperature, pressure, or speed; control modules monitoring actuator positions or valve states; and biological systems recording heart rate, oxygen saturation, or neural activity. Owing to their diverse origins and functions, each of these channels may differ in sampling resolution and noise characteristics, while also being interdependent, reflecting intricate inter-channel dependencies, where the behavior of one channel can influence or correlate with others. Given these realities, the temporal dynamics of each channel can vary significantly, and modeling all channels in the same manner can be both unrealistic and potentially misleading—especially when inter-channel dependencies are simply added without addressing channel-wise heterogeneity. Yet many existing approaches [34, 42, 40, 4] presuppose idealized conditions—such as uniform temporal granularity across channels and fully observed inputs at test time—that seldom hold in practical deployments. Sensor systems in complex environments often consist of diverse devices, each governed by distinct physical constraints, design specifications, and application requirements. Different channels may therefore be sampled at varying periods; for example, one sensor might collect data every hour, while another records measurements every 15 minutes. Consequently, even when aligned over the same overall time window, such discrepancies lead to input sequences of varying lengths across channels, complicating both model design and preprocessing workflows.
Figure 1: Newly proposed practical conditions include channel-wise asynchronous sampling periods and block-wise missing intervals at test time—further compound the long-standing challenge of inter-channel dependencies. Interpolation over coarsely sampled regions leads to signal distortion. See Appendix B for more details.
Despite these challenges, most standard time series forecasting models assume temporally aligned, fully observed inputs sampled at identical rates across channels. To conform to these assumptions, practitioners commonly resort to preprocessing techniques such as interpolation to impose regular intervals across channels, or downsampling, which reduces the resolution of high-frequency data to match the coarsest common sampling period across channels. While interpolation can introduce artificial continuity and spurious patterns, downsampling may cause loss of fine-grained temporal dynamics and eliminate informative fluctuations.
Beyond these structural discrepancies, real-world time series data are often further complicated by the presence of missing intervals at test time. In practice, sensors may malfunction temporarily or suffer communication failures, resulting in prolonged missing segments [48, 33], as opposed to sporadic individual gaps. Such contiguous absences pose a significant challenge in the input at test time, as accurate interpolation becomes difficult or even infeasible when missing intervals are long or the underlying signal is highly dynamic. In such cases, channel-dependent strategies offer a necessary alternative, enabling the model to infer missing dynamics from correlated signals across channels, rather than relying on unreliable interpolation.
These challenges point to a clear need for forecasting models that remain robust under more realistic conditions as illustrated in Figure 1. In particular, models should be capable of: (i) processing multivariate time series with asynchronous sampling and variable-length inputs across channels, without requiring alignment or resampling; (ii) handling block-wise missing intervals during test-time inference without strictly relying on prior imputation; and (iii) leveraging inter-channel correlations through channel-dependent strategies to recover missing dynamics and preserve structural consistency.
To tackle these challenges, we propose ChannelTokenFormer, a Transformer-based forecasting framework designed to operate under partial observability and asynchronous sampling across channels. The key innovation is the use of channel tokens—summary representations that aggregate information from local temporal tokens within each channel. These tokens act as global attention anchors, facilitating both intra-channel modeling and selective cross-channel information exchange. Moreover, only the channel tokens are passed to the decoder, enabling prediction directly from compressed channel-wise representations—regardless of how many local tokens each channel contains. This design allows the model to flexibly accommodate variable-length local representations across channels. Rather than enforcing strict alignment or relying on interpolation, the model incorporates frequency-adaptive patching and masking-aware attention mechanisms to preserve the natural resolution of each channel and operate robustly even under block-wise missing inputs. In summary, our key contributions are as follows.
  • We identify three practical but often overlooked challenges in time series forecasting: inter-channel dependencies, channel-wise asynchronous sampling, and block-wise test-time missing intervals.
  • We propose ChannelTokenFormer, a Transformer-based model that introduces channel tokens to capture channel-specific dynamics. It supports frequency-adaptive patching to handle asynchronous and variable-length inputs across channels without alignment. Furthermore, it performs unified attention with masking strategies that enable structured information exchange between local and channel-wise global tokens, model inter-channel dependencies for robust forecasting, and address block-wise missing values at test time.
  • Extensive experiments on multiple realistic benchmarks and a real-world dataset demonstrate the model’s effectiveness under input asynchrony caused by differing sampling periods, test-time missing intervals, and inter-channel dependencies—hallmarks of real-world time series.
2 Related Work
Multivariate Time Series Forecasting: Channel-Dependent Strategy
Channel-Dependent (CD) strategy treats all channels jointly to capture inter-channel dependencies, and has been widely adopted across various neural architectures, including TCN/CNNs [30, 44], MLPs [11, 39, 16], Transformers [6, 25, 35, 42, 40, 53, 29, 45, 49], and GNNs [15, 47, 1, 27]. These methods are particularly effective when channels are strongly correlated, as they enable the model to exploit complementary information across channels. In practical settings, it is common to encounter channel-wise differences in sampling periods and block-wise missing intervals at test time. Under such conditions, CD strategies can be particularly beneficial, as they allow sparsely observed or low-resolution channels to benefit from denser or more informative ones through cross-channel interactions. In this sense, channel dependence is not only useful but may in fact be essential for robust forecasting under non-ideal conditions.
Irregularly Sampled vs. Channel-wise Asynchronous Time Series
Irregularly sampled multivariate time series refer to data in which each channel is observed at non-uniform and often unpredictable time points, resulting in irregular gaps within individual channels. Several recent works [22, 5, 51, 31, 26, 13, 36, 52, 18, 24, 20] address this challenge by developing models that operate directly on non-uniform time intervals. These approaches often incorporate time encoding schemes, interpolation-based alignment, attention mechanisms, or graph-based structures within forecasting or classification tasks. While these methods address general forms of irregular sampling, our focus is on a more structured and practical setting in which each channel is sampled at a fixed period, but the periods differ across channels. In our setting, each channel is sampled at regular intervals, though the sampling periods differ across channels. This reflects real-world sensor systems, where each channel is recorded at its own period depending on the task requirements or hardware constraints. As a result, the input sequence becomes naturally asynchronous, with sparse observations that follow a clear and structured pattern—reflecting fixed yet distinct sampling periods across channels, in contrast to the random and irregular missingness assumed in many previous studies.
Missing Value Handling for Time Series and Test-time Missing Intervals
Early models such as BRITS [2] reconstruct missing values through imputation to reconstruct missing values based on bidirectional recurrent dynamics. General-purpose frameworks for time series, such as TimesNet [44] and TimeMixer++ [38], treat imputation as one of several downstream tasks. GRU-D [3] is designed for classification tasks on time series with missing values, while RAINDROP [52] extends this direction by explicitly modeling sensor-level missingness. However, in the context of forecasting tasks, the handling of test-time missing intervals—particularly block-wise gaps—remains underexplored. SERT [33] highlights the importance of handling block-wise missingness in real-world forecasting scenarios. While some studies [42, 56] evaluate model robustness under missing inputs, they typically replace missing values with zeros or mean values, without any explicit mechanism to address structured sparsity during test-time inference.
3 Problem Formulation
Asynchronous Channel-wise Observations
We consider a multivariate time series forecasting task where each channel $i\in\{1,\dots,N\}$ is sampled at a distinct, fixed period $s_{i}$. Let $s_{\min}=\min(s_{1},\dots,s_{N})$ denote the shortest sampling period, and define the sampling factor as $r_{i}=s_{i}/s_{\min}$. The input sequence for channel $i$ over a fixed observation window of length $L$ is represented as $\mathbf{x}^{(i)}=[x^{(i)}_{1},x^{(i)}_{2},\dots,x^{(i)}_{L_{i}}]\in\mathbb{R}^{L_{i}}$, where $L_{i}=\lfloor L/r_{i}\rfloor$. Given a prediction window of length $H$, the forecast horizon for channel $i$ becomes $\mathbf{y}^{(i)}=[y^{(i)}_{1},y^{(i)}_{2},\dots,y^{(i)}_{H_{i}}]\in\mathbb{R}^{H_{i}}$, where $H_{i}=\lfloor H/r_{i}\rfloor$. This implies that channels with longer sampling periods contain fewer observations within the same time window, resulting in shorter prediction horizons $H_{i}$. Note that this adjustment is made in the data point domain, not the time domain, so that the actual time horizon covered by each channel’s prediction remains consistent. We formulate a forecasting model $f_{\theta}$ parameterized by $\theta$ that predicts the future values by leveraging sampling period information from all channels: $\hat{\mathbf{y}}=f_{\theta}(\{(\mathbf{x}^{(i)},s_{i})\}_{i=1}^{N})$. Here, $s_{i}$ enables the model to account for the temporal resolution and data density of each input channel. The learning objective is to minimize the mean squared error across all channels, each of which may have a different prediction length $H_{i}$ due to its unique sampling period. The total training loss is computed by first calculating the average MSE over each channel’s prediction horizon, then averaging across channels. We refer to this as Channel-aggregated MSE (CMSE):
$\mathcal{L}_{\text{total}}=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}^{(i)}=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{H_{i}}\frac{1}{H_{i}}\left(y^{(i)}_{j}-\hat{y}^{(i)}_{j}\right)^{2}.$ (1)
For evaluation, we adopt CMSE as the primary metric, along with CMAE, its MAE-based variant.
Test-time Missing Intervals
In many real-world scenarios, observations exhibit persistent sparsity due to temporary failures, communication losses, or maintenance, leading to block-wise missing intervals at test time. This type of missingness is typically easy to detect, as most channels rarely show prolonged zero readings unless the device is disconnected or malfunctioning. Accordingly, we introduce a binary mask as auxiliary metadata to indicate whether each channel contains valid observations:
$\mathbf{m}^{(i)}\in\{0,1\}^{L_{i}},\quad m^{(i)}_{j}=\begin{cases}1,&\text{if }x_{j}^{(i)}\text{ is observed},\\ 0,&\text{if }x_{j}^{(i)}\text{ is missing}\end{cases}.$ (2)
The binary mask $\mathbf{m}^{(i)}$ indicates whether each patch in channel i should be masked or retained. Conventional patch-based methods cannot handle variable numbers of patches per channel, and thus are unable to leverage this binary mask. Instead, they typically fill missing entries within each patch using zeros or mean values. In contrast, our model adopts a non-overlapping patch representation that supports variable patch lengths and counts across channels. This enables the binary mask to be effectively used for removing fully unobserved patches from the input, rather than merely filling in missing values. Accordingly, the forecasting model is extended to accommodate masking-aware inputs across all channels: $\hat{\mathbf{y}}=f_{\theta}(\{(\mathbf{x}^{(i)},\mathbf{m}^{(i)},s_{i})\}_{i=1}^{N})$. This formulation supports forecasting under heterogeneous horizons from channel-wise asynchronous sampling, while also accommodating channel-specific missingness in the input at test time. The forecast horizon is assumed to be fully observed, since missing intervals in the future cannot be predicted in advance.
4 ChannelTokenFormer
Figure 2: Overview of ChannelTokenFormer (CTF). All tokens across channels pass through a unified attention layer, where local and global information is aggregated into channel tokens. Only the channel tokens are decoded by a decoder shared per sampling period to produce the final prediction.
Motivation for Channel Tokens
While some multivariate time series are regularly sampled, many real-world scenarios involve channel-specific sampling periods and prolonged test-time missing intervals. Such conditions limit the applicability of forecasting models that assume fixed-length, fully observed, and temporally aligned inputs. Although pre-filling is commonly used, it often distorts signal characteristics and induces artificial temporal patterns, ultimately degrading forecasting accuracy under asynchronous sampling and missingness.
To address these limitations, we propose Channel Tokens—learned global representations that summarize temporal dynamics within each channel and support forecasting under asynchronous sampling, test-time missingness, and inter-channel dependencies. Channel tokens attend to the observed patch-level inputs of their corresponding channels via masked self-attention, eliminating the need for aligned or interpolated sequences. This design enables the model to handle asynchronous sampling and missing intervals by compressing local information into channel tokens and enabling cross-channel interaction through attention among them.
Channel-wise Frequency-based Dynamic Patching and Tokenization
Each input channel is represented as a one-dimensional sequence $\mathbf{x}^{(i)}$. To determine an appropriate patch length, we apply the Fast Fourier Transform (FFT) to obtain the amplitude spectrum $\mathbf{A}^{(i)}=|\hat{\mathbf{x}}^{(i)}|$. For channel $i$, if a strong dominant frequency $f^{*}_{i}=\arg\max_{k}\mathbf{A}^{(i)}_{k}$ is detected, we set the patch length to its corresponding period: $\ell_{i}=\left\lfloor L_{i}/f^{*}_{i}\right\rfloor$. Otherwise, the patch length is determined based on the channel’s relative sampling period. This frequency-aware patching allows each channel to be divided into non-overlapping local patches that reflect its temporal characteristics, which are later summarized by channel tokens. Each local patch $\mathbf{p}^{(i)}_{j}\in\mathbb{R}^{\ell_{i}}$ is projected into a fixed-dimensional token $\mathbf{L}^{(i)}_{j}\in\mathbb{R}^{d}$ using a linear layer specific to its patch length: $\mathbf{L}^{(i)}_{j}=\mathbf{W}_{\text{patch}}^{(\ell_{i})}\mathbf{p}^{(i)}_{j}$. To improve efficiency, patches with the same length share a projection layer, enabling tokenizers to be reused across channels. Because patch counts and lengths vary across channels, this design balances efficiency and flexibility. In parallel, multiple channel tokens $\mathbf{C}^{(i)}=[\mathbf{C}^{(i)}_{1};\dots;\mathbf{C}^{(i)}_{C}]\in\mathbb{R}^{C\times d}$ are initialized for each channel to serve as global representations, where $C$ denotes the number of channel tokens for each channel. We add a fixed positional embedding $\mathbf{e}_{\text{pos}}$ and a learnable channel embedding $\mathbf{e}_{\text{ch}}^{(i)}$, which indicates the identity of channel $i$, to each local token, while the channel token receives only the channel embedding.
Input-Length Variability through Channel-wise Patch Masking
Building on the flexibility of channel-wise dynamic patching, our framework allows for variable input lengths via patch masking. Leveraging this capability, we propose a patch masking strategy to handle interval missingness at test time. Specifically, if a missing segment spans an entire patch identified by the binary mask, the corresponding local token is excluded from the attention computation. As illustrated in Figure 2, this prevents fully unobserved patches from contributing to attention, thereby reducing the risk of propagating invalid information. Despite temporal gaps from masking, the model retains and understands temporal context through fixed positional embeddings. To improve robustness, we propose randomly masking a subset of channel-wise patches during training [28, 34]. This strategy improves resilience to test-time missing intervals and also serves as an implicit regularizer, reducing overfitting in low-resource forecasting scenarios.
Unified Attention via Masked Local and Channel Tokens
To jointly model temporal patterns within each channel and cross-channel dependencies, we design a unified masked self-attention mechanism that integrates both local and global representations into a single attention operation (see Figure 2). Each channel $i\in\{1,\dots,N\}$ contributes two types of tokens: (i) local tokens $\mathbf{L}^{(i)}\in\mathbb{R}^{L^{(i)}\times d}$ representing patch-level embeddings, and (ii) global channel tokens $\mathbf{C}^{(i)}\in\mathbb{R}^{C\times d}$ summarizing high-level contextual information, where $L^{(i)}$ and $C$ denote the number of local and channel tokens, respectively. The full token sequence is constructed as $\mathbf{X}=[\mathbf{L}^{(1)};\mathbf{C}^{(1)};\dots;\mathbf{L}^{(N)};\mathbf{C}^{(N)}]\in\mathbb{R}^{T\times d}$, with total token count $T=\sum_{i=1}^{N}(L^{(i)}+C)$. This unified sequence is passed through a masked multi-head self-attention layer with residual connection: $\mathbf{X}_{\text{out}}=\mathbf{X}+\text{Attention}(Q,K,V)$, where $Q=\mathbf{X}\mathbf{W}_{Q}$, $K=\mathbf{X}\mathbf{W}_{K}$, and $V=\mathbf{X}\mathbf{W}_{V}$ are linear projections of $\mathbf{X}$, and the attention is computed as $\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d}}+\mathbf{M}\right)V$. Here, $\mathbf{M}\in\mathbb{R}^{T\times T}$ encodes structural constraints on how tokens can attend to one another, reflecting their types, channels, and masking. Our tailored masking scheme defines the construction of $\mathbf{M}$ as follows (See Figure 3): (1) Local tokens attend only to other local tokens within the same channel, enabling intra-temporal modeling. (2) Channel tokens attend to their own local tokens and to other channels’ tokens, but are not accessible to local tokens due to their read-only role in the attention mechanism. (3) Channel tokens do not attend to themselves; for each query $\mathbf{C}^{(i)}$, the key $\mathbf{C}^{(i)}$ is excluded to avoid self-reinforcement and to encourage informative cross-channel interaction. Our unified masking strategy, inspired by read-only prompt attention [21] and attention control techniques [19], is implemented without modifying the standard Transformer architecture. Stacked masked attention blocks with residual connections and layer normalization enable deeper modeling of temporal and cross-channel dependencies.
While the channel-wise design of CTF can lead to parameter overhead, unifying local and cross-channel attention into a single layer helps offset this cost and improves parameter efficiency. This design provides an inductive bias that simultaneously supports temporal abstraction, captures channel dependencies, and enhances robustness to missing intervals and asynchronous sampling—key challenges in real-world forecasting.
Figure 3: Our unified attention masking strategy. Local tokens perform intra-temporal attention within the same channel. Channel tokens aggregate local and cross-channel information, but are not visible to local tokens and do not attend to themselves. Optionally, attention among channel tokens from the same channel can be masked to encourage inter-channel interaction and reduce redundancy.
5 Experiments
To assess the effectiveness of ChannelTokenFormer under the newly proposed problem settings, we conduct extensive experiments on three multivariate time series benchmark datasets that we modify to reflect practical conditions, and on one real-world dataset from the LNG cargo handling system of an LNG carrier.
Datasets
We evaluate forecasting performance under channel-wise asynchronous sampling using four multivariate datasets. ETT1-practical (ETT1) modifies ETTm1 [55] by resampling channels to domain-specific temporal resolutions (e.g., 1-hour for load, 15-minute for temperature). Weather-practical (Weather), adapted from the Weather benchmark [43], assigns heterogeneous sampling periods to channels based on distinct physical characteristics. Monash-SolarWindPower-practical (SolarWind) combines solar and wind power series from the Monash Forecasting Archive [10], with variables resampled to 20-minute and 5-minute intervals, respectively. LNG Cargo Handling System (CHS) is a real-world industrial dataset collected from an LNG carrier, consisting of sensor channels related to cargo operations, ship navigation, and surrounding weather conditions. To assess robustness under block-wise test-time missing intervals, we additionally conduct experiments on SolarWind by varying the missing ratio. Detailed dataset specifications are provided in Appendix A.1.
Baselines
To provide a comprehensive evaluation, we compare our CTF against representative state-of-the-art models from diverse architectural paradigms. We include Transformer-based models such as TimeXer [42], iTransformer [29], and PatchTST [34], along with the CNN-based TimesNet [44], the GNN-based CrossGNN [15], and MLP-based models TimeMixer [39] and DLinear [50].
Setup and Implementation Details
Most baseline models assume regularly sampled and fully observed inputs. To meet this assumption, we use linear interpolation or forward filling to recover values unobserved due to differences in channel-specific sampling periods. However, to simulate test-time block-wise missingness, we randomly select contiguous regions in the input to be masked and apply zero-filling instead of other interpolation methods. This decouples the impact of missingness from asynchronous sampling and creates a conservative evaluation setting that amplifies differences between models with and without explicit missing-aware mechanisms. For each dataset, input and prediction lengths are selected to reflect the channel-wise temporal resolution. To accommodate heterogeneous sampling periods across channels, we define these lengths based on a fixed time duration rather than a fixed number of steps. The actual number of input and output points per channel is determined by its sampling period, with shorter periods yielding more points over the same window. Detailed settings are provided in Appendix A.2.
5.1 Main Results
Scenario 1: Channel-wise Asynchronous Forecasting
We evaluate forecasting performance without test-time missing intervals using our newly constructed datasets that reflect realistic sampling heterogeneity across channels. Unlike baselines that rely on interpolation, our approach preserves the original sampling structure. As summarized by average CMSE and CMAE for all prediction lengths in Table 1, our method consistently outperforms all baselines across datasets in the practical setting. These results highlight the advantage of avoiding interpolation in asynchronous environments. Detailed results for each prediction length and dataset are provided in Appendix E.
Table 1: (Scenario 1) Forecasting performance on the channel-wise asynchronous forecasting task.
Strategy
Channel-Dependent
Channel-Independent
Models
CTF(ours)
TimeXer
TimeMixer
iTrans.
CrossGNN
TimesNet
PatchTST
DLinear
Metric
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
ETT1
0.398
0.411
0.420
0.424
0.420
0.417
0.434
0.432
0.430
0.418
0.468
0.449
0.414
0.418
0.424
0.419
SolarWind
0.400
0.448
0.420
0.468
0.449
0.486
0.463
0.481
0.469
0.487
0.477
0.478
0.412
0.463
0.420
0.514
Weather
0.269
0.293
0.297
0.308
0.295
0.306
0.312
0.322
0.284
0.313
0.311
0.322
0.302
0.317
0.286
0.330
CHS
0.282
0.126
0.297
0.128
0.299
0.127
0.308
0.132
0.286
0.141
0.334
0.141
0.316
0.132
0.338
0.236
Scenario 2: Channel-wise Asynchronous Forecasting with Block-wise Test-time Missing
We further evaluate all models under a more challenging setting, where block-wise missing intervals are applied to test-time inputs, on top of channel-wise asynchronous sampling. In this scenario, all models receive inputs containing contiguous missing segments filled with zeros. Although zero-filling is applied for input compatibility, our model uniquely utilizes masking to prevent attention to regions treated as missing. As shown in Table 2, our model consistently outperforms all baselines in average CMSE and CMAE for all prediction lengths across different missing ratios, maintaining top performance even as the severity of missingness increases. Notably, at a missing ratio of 0.25, the performance degradation of our model remains as low as 1.23%, indicating its strong resilience to partially missing inputs. This robustness stems from the use of random patch masking during training, which improves the model’s ability to generalize to incomplete patterns. Unlike other methods, our model avoids distortion from artificially zero-filled values and preserves signal fidelity under block-wise missing conditions. Detailed results across datasets are provided in Appendix E.
Table 2: (Scenario 2) Forecasting performance on the SolarWind dataset under block-wise test-time missingness with varying missing ratios $m$.
Strategy
Channel-Dependent
Channel-Independent
Models
CTF(ours)
TimeXer
TimeMixer
iTrans.
CrossGNN
TimesNet
PatchTST
DLinear
Metric
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
$m=0.0$
0.405
0.455
0.420
0.468
0.449
0.486
0.463
0.481
0.469
0.487
0.477
0.478
0.412
0.463
0.420
0.514
$m=0.125$
0.396
0.458
0.432
0.491
0.445
0.499
0.460
0.490
0.467
0.509
0.465
0.485
0.452
0.515
0.432
0.534
$m=0.250$
0.410
0.477
0.473
0.521
0.464
0.522
0.479
0.513
0.474
0.535
0.482
0.511
0.527
0.585
0.450
0.556
$m=0.375$
0.424
0.500
0.507
0.544
0.499
0.550
0.507
0.540
0.490
0.562
0.502
0.543
0.600
0.638
0.472
0.579
$m=0.500$
0.452
0.530
0.543
0.564
0.540
0.575
0.556
0.570
0.510
0.591
0.533
0.577
0.663
0.668
0.496
0.602
5.2 Ablation Study
Table 3: Ablation study under channel-wise asynchronous sampling with block-wise test-time missing (missing ratio = 0.25). Results are averaged across all prediction lengths on the SolarWind dataset.
Ablation Setting CMSE CMAE
[…]
w/o Channel Dependence (CI only) 0.435 0.484
w/o Dynamic patching 0.430 0.492
w/o Patch masking 0.486 0.525
To assess the effectiveness of our design in addressing the three key challenges—Dependency, Asynchrony, and Missingness—we conduct a structured ablation study in which each component of our architecture is selectively removed or replaced. For Dependency, we disable cross-channel attention by restricting interactions between channel tokens from different channels. In the case of Asynchrony, we remove the dynamic patching mechanism, forcing the model to process input sequences with varying lengths using fixed patch boundaries that ignore dominant frequencies and channel-specific sampling periods. For Missingness, we eliminate patch masking during both training and testing, thereby exposing the model to inputs with artificially zero-filled values in place of structured missingness handling. Robust forecasting in practical settings requires all three components—channel dependence, dynamic patching, and patch masking—as the removal of any one leads to degraded performance. Additional ablations are provided in Appendix C.1.
6 Analysis
6.1 Effect of Number of Channel Tokens
To examine the impact of channel token granularity, we vary the number of channel tokens across all practical datasets. As shown in Table 5, the optimal channel token count differs by dataset, reflecting variations in size, temporal complexity, and signal diversity. Two tokens per channel yield the best performance on ETT1 and SolarWind, balancing expressiveness and generalization. Weather benefits from more tokens due to its multi-scale, high-variability signals, whereas CHS performs best with a single token, due to sparse, low-frequency control signals where over-tokenization leads to overfitting. These results suggest tuning the number of channel tokens based on dataset characteristics—favoring fewer tokens for small or redundant datasets, and more for rich, dynamic ones.
Table 4: Effect of varying the number of channel tokens {1, 2, 3} on asynchronous forecasting performance. Results are reported in CMSE, averaged over all prediction lengths.
# Channel Tokens ETT1 SolarWind Weather CHS
[…]
Table 5: Evaluation of robustness to varying input lengths during test time. Our CTF is trained with an input length of 576 using random patch masking, and tested with other shorter input lengths of 288, 360, and 432 on the SolarWind dataset.
Test Input Length 288 360 432 576
Avg. CMSE 0.429 0.461 0.448 0.409
Avg. CMAE 0.457 0.495 0.491 0.457
6.2 Robustness to Input Length Variability
Unlike conventional forecasting models restricted to fixed-length inputs, our method is trained with random patch masking to improve robustness to variable-length sequences. As shown in Table 5, the model maintains stable performance across all test input lengths on the SolarWind dataset without requiring retraining. Since each channel has a different sampling period, we define the input length based on the channel with the smallest sampling interval, which corresponds to 576 time steps over 2 days. Even when this length is reduced to 288 (equivalent to 1 day), the CMSE increases by only 4.9%, demonstrating strong generalization to input length variation, which is an essential property for real-world forecasting under inconsistent input conditions.
6.3 Frequency Bias Analysis in Predicted Time Series
Figure 4: Frequency-domain comparison between CTF and TimeXer on a test sample from ETT1, showing that amplitude attenuation is prevented across all frequency bands.
We evaluate the impact of interpolation on the frequency characteristics of predicted time series under practical conditions with channel-wise asynchronous sampling periods. To this end, we compare the spectral fidelity of forecasts from our interpolation-free model, ChannelTokenFormer (CTF), with that of TimeXer, a representative interpolation-based baseline. The analysis focuses on dominant frequency difference and amplitude spectrum RMSE computed across low, mid, and high frequency bands, with Table 6 reporting overall means averaged across all channels in the ETT1 dataset. Zero-centering is applied prior to spectral analysis to remove ultra-low-frequency (DC) components and enhance periodicity detection. In the ETT1 dataset, dominant periodicities are concentrated in the low- to mid-frequency range (e.g., 8, 12, and 24 hours), where interpolation-induced phase delay has limited effect. As a result, even TimeXer, which relies on interpolation, is able to localize the dominant spectral peak with an accuracy comparable to that of CTF. However, the amplitude spectrum RMSE reveals consistent discrepancies across all frequency bands, indicating that CTF more effectively preserves the magnitude of spectral components. As illustrated in Figure 4, TimeXer undergoes greater amplitude attenuation compared to CTF. In contrast, CTF better maintains the original spectral energy. These results indicate that CTF mitigates spectral distortion introduced by interpolation and preserves frequency structure under asynchronous sampling, which translates to improved trend and periodicity retention in the time domain.
Table 6: Frequency-domain comparison between CTF (interpolation-free) and TimeXer (interpolation-based) on the ETT1 dataset for a prediction length of 192 (2-day window). Lower values indicate better spectral fidelity. RMSEs are computed over three frequency bands under 1-hour sampling ($f_{s}=1$): Low (0.00–0.10), Mid (0.10–0.25), and High (0.25–0.50) cycles/hour, corresponding to periodicities longer than 10 hours, between 4 and 10 hours, and shorter than 4 hours, respectively.
Model / Metric Dominant Freq. Diff. Low-band RMSE Mid-band RMSE High-band RMSE
[…]
TimeXer 0.0099 0.1901 0.0745 0.0473
7 Conclusion
We introduce a practical forecasting setting that reflects real-world conditions, including channel-wise asynchronous sampling and block-wise missing intervals. Instead of relying on interpolation, our approach handles both missing segments and sampling gaps through masking and frequency-based dynamic patching. Built on this setting, our unified attention model leverages channel tokens with read-only masking and cross-channel attention, consistently outperforming prior methods across real-world scenarios. However, as the number of channels increases, full attention across all channels, based purely on masking to regulate information flow, may scale less efficiently. To support high-dimensional forecasting, future work will focus on developing more scalable versions of our model.
Appendix
Appendix A Implementation Details
A.1 Dataset Configuration under Channel-wise Asynchrony
We conduct long-term forecasting experiments on three modified benchmark datasets adapted to our practical setting, as well as one real-world dataset. In this setting, each channel provides a different number of observations over the same temporal window, reflecting channel-specific sampling periods in the multivariate time series. This configuration more closely mirrors real-world conditions, where sensor deployment strategies and data acquisition policies vary across channels.
ETT1-practical (ETT1)
ETT1-practical is a variant of the commonly used ETTm1 [55] dataset. Unlike the original version, where all channels are sampled uniformly at 15-minute intervals, ETT1-practical introduces channel-specific sampling periods to reflect industrial environments, where sensors operate at different acquisition frequencies depending on their physical properties and monitoring requirements. Specifically, the channels are grouped into two categories: load-related channels and the oil temperature (OT) channel. The load-related channels are downsampled to an hourly rate, representing slowly varying electrical load measurements, while the OT channel retains its original 15-minute resolution to capture high-frequency thermal dynamics.
Weather-practical (Weather)
Weather-practical is constructed from the widely used Weather [43] benchmark. While the original dataset assumes uniform sampling periods across all channels, this assumption rarely holds in real-world environmental monitoring systems. In Weather-practical, we reorganize the channels to reflect heterogeneous sampling frequencies, based on the intrinsic temporal dynamics of each channel. For instance, fast-changing environmental channels such as wind velocity, precipitation, and solar radiation are sampled at 10-minute intervals, whereas more stable channels such as air pressure or vapor pressure are sampled at a coarser hourly rate. Mid-range channels—such as air temperature and specific humidity—are sampled at 20-minute or 30-minute intervals in accordance with their moderate temporal variability. This configuration, summarized in Table 7, mimics the asynchronous and heterogeneous acquisition patterns encountered in operational weather stations and provides a more realistic testbed for evaluating forecasting models under non-uniform input conditions.
Table 7: Channel Specification in the Weather-practical Dataset
Channel Description Sampling Period
p (mbar) Air Pressure 1 hour
T (C) Air Temperature 20 min
Tpot (K) Potential Temperature 1 hour
Tdew (C) Dew Point Temperature 1 hour
rh (%) Relative Humidity 30 min
VPmax (mbar) Saturation Water Vapor Pressure 1 hour
VPact (mbar) Actual Water Vapor Pressure 1 hour
VPdef (mbar) Water Vapor Pressure Deficit 1 hour
sh (g/kg) Specific Humidity 1 hour
H2OC (mmol/mol) Water Vapor Concentration 1 hour
rho (kg/m3) Air Density 1 hour
wv (m/s) Wind Velocity 10 min
max. wv (m/s) Max Wind Velocity 10 min
wd () Wind Direction 10 min
rain (mm) Precipitation 10 min
raining (s) Duration of Precipitation 10 min
SWDR (W/m2) Short Wave Downward Radiation 10 min
PAR ($\mu$mol/m2/s) Photosynthetically Active Radiation 10 min
max. PAR ($\mu$mol/m2/s) Max PAR 10 min
Tlog (C) Internal Logger Temperature 1 hour
OT Operational Timestamp (Offset) 10 min
Monash-SolarWind-practical (SolarWind)
Monash-SolarWind-practical is derived from the Monash Forecasting Repository [10] and built upon high-frequency power generation data originally collected at 4-second intervals. In real-world energy monitoring systems, distinct sensors are used for solar and wind power, each exhibiting unique temporal behaviors and operational constraints. To replicate these real-world dynamics, we apply channel-specific downsampling: solar-related channels are resampled at 20-minute intervals, capturing smooth and gradual irradiance trends, while wind-related channels are resampled at 5-minute intervals to reflect the rapid fluctuations characteristic of wind dynamics. This restructured dataset captures modality-specific asynchrony and sampling heterogeneity, offering a challenging benchmark for assessing model robustness under realistic conditions where different signal types operate on distinct temporal resolutions.
LNG Cargo Handling System (CHS)
CHS is a real-world dataset collected from an LNG (Liquefied Natural Gas) carrier during ballast voyage. The dataset contains time series measurements from a total of 52 channels, each sampled at distinct frequencies based on their physical characteristics and operational requirements. Specifically, weather and navigation channels, such as latitude and wave period, are sourced from external systems and sampled at a coarse hourly rate. In contrast, onboard sensor channels are sampled more frequently to support real-time monitoring and control. These include machinery-related channels (e.g., Low Duty Compressor (LDC) suction pressure, main engine inlet pressure, and LNG consumption), sampled at 10-minute intervals, and cargo tank sensor channels (e.g., tank temperature at top, mid, and bottom levels, and tank pressure), sampled at 1-minute intervals. For experimental evaluation, we selected 10 key channels from the full set of 52, focusing on those most relevant to navigation and weather conditions, cargo tank thermodynamics, and machinery operations.
The LNG carrier is equipped with four independent cargo tanks. During voyage, ambient heat ingress causes the LNG stored in these tanks to partially vaporize, producing what is known as Boil-Off Gas (BOG). Efficient BOG management is critical: the gas can be used as fuel, reliquefied, or combusted via a Gas Combustion Unit (GCU), which converts it into CO2 before discharge. The operational objective is to minimize unnecessary combustion while maximizing energy utilization as fuel, all while maintaining tank pressure within safe operational bounds. The CHS dataset captures the dynamic interplay among external conditions, cargo tank behavior, and onboard machinery responses. It provides a foundation for predictive modeling of key control channels, including tank gas pressure and LDC suction pressure. Accurate prediction of these channels is important for safe and energy-efficient voyage execution under the channel-wise asynchronous sampling conditions found in real-world LNG cargo operations.
Figure 5: Overview of the LNG Cargo Handling System (CHS).
A.2 Implementation Details
Model Hyperparameters and GPU Setup
All experiments were conducted using a single NVIDIA RTX 3090 GPU with 24GB memory. We used the Adam optimizer across all models, with initial learning rates per dataset for fair comparison: $10^{-4}$ for ETT1 and SolarWind, $10^{-2}$ for Weather, and $10^{-3}$ for the CHS dataset. Training was run for a maximum of 10 epochs with early stopping based on validation performance. For models with a stackable encoder block, the number of encoder layers was set to 2. For Transformer-based models, the model dimension $d_{\text{model}}$ was searched over $\{128,256,512\}$, and the feedforward expansion ratio $d_{\text{ff}}$ / $d_{\text{model}}$ was selected from $\{1,2,4\}$. We adopted most baseline implementations from the TimesNet [44] repository. CrossGNN [15], originally provided in a separate repository, was reimplemented and integrated into the TimesNet framework for consistency. In our proposed ChannelTokenFormer (CTF), the number of channel tokens was selected from $\{1,2,3\}$, with the optimal value determined separately for each dataset.
Training and Evaluation under Our Proposed Practical Setting
For baseline models, both training and testing were conducted on linearly interpolated data due to architectural constraints. To ensure a fair and consistent comparison, we preserved the original design of the baselines and trained them on interpolated inputs with MSE loss. For additional comparison, we also evaluated baseline models under the practical setting using forward-filled inputs. Empirically, we found that their performance remained largely consistent with that obtained using linear interpolation, suggesting that the interpolation method had limited influence on overall accuracy for these models (See Table 12).
In contrast, our proposed model, ChannelTokenFormer (CTF), is explicitly designed to operate on non-interpolated inputs that reflect the original, variable sampling periods of each channel. To ensure uniform input shape compatible with existing architectures, we apply forward-fill interpolation during dataset preprocessing; however, this step serves merely as a structural tool to meet input dimensionality requirements, not as a modeling assumption. During the channel-wise patch embedding phase, all forward-filled values are explicitly excluded: only the valid indices $L_{i}$, corresponding to truly observed data points, are used for computation.
To ensure fair evaluation, both the baselines and our model compute channel-aggregated metrics—CMSE and CMAE, as defined in Section 3—only over the valid indices $H_{i}$, which correspond to actually observed target values. Thus, although the input preprocessing methods differ, all models are evaluated on the same set of ground truth targets $(y_{1},\dots,y_{H_{i}})$. Note that under the practical setting, the number of valid test samples is inherently reduced, since the test set is subsampled according to the maximum sampling factor across channels. In summary, each model is trained under settings aligned with its architectural requirements, but all are evaluated uniformly based on their accuracy in predicting truly observed values.
Input/Output Length Settings per Dataset
As described in Section 5, we set the input and prediction lengths for each dataset to reflect the temporal characteristics of the underlying signals. Instead of fixing the number of time steps across all channels, we define these lengths over a consistent time duration to accommodate channel-wise variations in both sampling period and dominant frequency components. For implementation and reporting convenience, however, we represent input and output lengths in terms of the number of steps corresponding to the channel with the smallest sampling period. The input and prediction lengths (in time steps, relative to the finest resolution channel) for each dataset are as follows: ETT1: 192 (input) $\rightarrow$ [192, 336, 768, 1152] (prediction), SolarWind: 576 $\rightarrow$ [288, 576, 864, 1152], Weather: 144 $\rightarrow$ [144, 288, 576, 864], CHS: 120 $\rightarrow$ [120, 240, 480, 840].
Example of Channel-wise Dynamic Patching
In our model, the number of local tokens per channel adapts to the characteristics of its signal: channels with finer temporal resolution or shorter dominant periods, as determined via FFT, result in more tokens, whereas coarser or slower signals result in fewer tokens within the same time window. For example, in the SolarWind dataset, wind power is recorded every 5 minutes and solar power every 20 minutes. Over a 2-day period, this corresponds to 576 time steps for wind power (dominant period: 48 steps) and 144 for solar power (dominant period: 18 steps). Assuming patching is performed based on each channel’s dominant period, this yields 576 / 48 = 12 local tokens for wind power and 144 / 18 = 8 for solar power.
Test-time Missing Conditions
To evaluate robustness under structured missing conditions, we introduce block-wise missing intervals in the test-time inputs of two datasets. In the SolarWind dataset, each channel is assigned randomly positioned missing blocks, with each block spanning exactly one patch length. This ensures that at least one patch per channel is entirely unobserved, enabling a direct assessment of our model’s patch-level masking strategy at test time. During training, our model applies random patch masking, as shown in Table 15, training with a 0.4 masking ratio consistently improved model robustness across different missing ratios. In the ETT1 dataset, we insert random zero-filled missing blocks of varying lengths (5–20 time steps) into each channel. These settings are designed to test model robustness to short-range missing intervals. Detailed results under this setting are reported in Table 16.
Appendix B Spectral Distortion Induced by Interpolation
B.1 Background
In practical time series forecasting scenarios, multivariate signals often feature channel-wise asynchronous sampling periods across channels, stemming from both inter-system and intra-system heterogeneity in sensor design and operation. However, most existing forecasting models assume a regular and complete time grid, necessitating the use of interpolation to align all channels to a common temporal resolution prior to model ingestion. While this preprocessing step ensures input compatibility, it inevitably introduces spectral distortion [8, 9] by artificially synthesizing intermediate values, thereby altering the original frequency characteristics of the signals. Interpolation imposes artificial continuity by estimating unobserved values from nearby observations, which can fundamentally reshape both the temporal dynamics and spectral content of the signal. This appendix provides a theoretical explanation of how such interpolation-induced distortion arises in the frequency domain, as observed through the Fast Fourier Transform (FFT).
B.2 Linear Interpolation-induced Frequency Domain Distortion
Let $x(t)$ denote a real-valued time series obtained via uniform but sparse sampling relative to other channels. Let $\tilde{x}(t)$ denote the version of $x(t)$ with intermediate values filled in via linear interpolation between uniformly sampled points. This interpolation process can be interpreted as a convolution in the time domain, as linear interpolation is equivalent to applying a piecewise linear (triangular) filter to the sampled signal. By the Convolution Theorem, this results in a multiplication in the frequency domain:
$\tilde{X}(f)=X(f)\cdot H(f),$
where $H(f)$ is the frequency response of the triangular interpolation kernel, given by $\mathrm{sinc}^{2}(f)$. This imposes a low-pass filtering effect that attenuates high-frequency components of the original signal, and constitutes the primary source of spectral distortion introduced by linear interpolation. Such distortion manifests in multiple forms: (i) amplitude attenuation within the Nyquist limit, due to low-pass filtering that suppresses mid-to-high frequency energy; (ii) amplitude gaps, due to the non-uniform and oscillatory nature of the interpolation kernel’s frequency response, which causes broad spectral suppression in the mid-to-high frequency range; (iii) spectral leakage beyond the Nyquist frequency, caused by the non-ideal frequency response of finite-support interpolation kernels; and (iv) phase delays, introduced by the filtering process, especially in the mid-to-high frequency range, and observable in the FFT phase spectrum.
(1) Amplitude Attenuation
Under linear interpolation, the reconstructed signal exhibits reduced overall spectral energy, particularly in the mid-to-high frequency range:
$\|\tilde{X}(f)\|_{2}^{2}<\|X(f)\|_{2}^{2}.$
This occurs because each frequency component is scaled by a factor $|H(f)|<1$ for $f>0$, resulting in suppressed spectral magnitude:
$|\tilde{X}(f)|=|X(f)|\cdot|H(f)|<|X(f)|.$
Attenuation becomes stronger at higher frequencies, leading to a net reduction in total spectral energy and a smoother appearance in the time domain. This loss of high-frequency information can impair the model’s ability to capture fine-grained temporal patterns in forecasting tasks.
(2) Amplitude Gaps
The frequency response of linear interpolation exhibits oscillatory behavior due to its squared sinc form, resulting in repeated dips and low-gain regions across the spectrum, particularly in the mid-to-high frequency range. When these regions coincide with frequencies that contain substantial energy in the original signal, the interpolated spectrum displays significant local suppression—observable as amplitude gaps:
$|\tilde{X}(f)|\ll|X(f)|\quad\text{for }f\in[f_{\text{mid}},f_{\text{high}}].$
Such non-uniform attenuation alters the spectral envelope and can adversely impact the performance of forecasting models.
(3) Spectral Leakage (High-Frequency Artifacts)
The frequency response of linear interpolation, given by $H(f)=\mathrm{sinc}^{2}(f)$, is not band-limited and extends well beyond the Nyquist frequency, with decaying side lobes across the spectrum. As a result, interpolation artificially introduces high-frequency components into the signal, even when the original data is strictly band-limited. This is particularly evident in the FFT magnitude spectrum(see Figure 6(a)), where the interpolated signal exhibits spectral energy in regions beyond the Nyquist limit. These high-frequency artifacts are not part of the true signal but emerge from the spectral spreading effect caused by the interpolation kernel. They distort the original spectral structure by injecting spurious out-of-band energy, which can mislead models.
(4) Phase Delay
Although linear interpolation produces a smooth and continuous waveform from discretely sampled data, it introduces a frequency-dependent shift in the phase spectrum. Let $\angle X(f)$ and $\angle\tilde{X}(f)$ denote the phase spectra of the original and interpolated signals, respectively. Then the phase difference is given by:
$\Delta\phi(f)=\angle\tilde{X}(f)-\angle X(f)\leq 0,$
which remains negative across most frequencies, with the delay becoming notably apparent from the mid-frequency band onward (see Figure 6(b)). This indicates that the interpolated signal lags behind the original in phase. The delay is negligible at low frequencies but grows rapidly in the high-frequency band, due to the squared sinc frequency response of linear interpolation. When the signal’s dominant spectral components lie in the high-frequency region, this phase delay leads to substantial misalignment of key waveform features such as peaks and transitions. Even if the interpolated waveform appears visually smooth in the time domain, the underlying temporal displacement of high-frequency content can significantly degrade forecasting accuracy, especially in tasks that rely on precise timing of local patterns.
B.3 Summary and Implication
The four types of distortion—amplitude attenuation, amplitude gaps, spectral leakage, and phase delay—collectively degrade the spectral integrity of interpolated signals. These artifacts can mislead forecasting models into learning spurious temporal or frequency patterns. The issue becomes more severe in multivariate settings with channel-wise asynchronous sampling, where aligning all channels to a uniform time grid via interpolation introduces artificial synchrony and continuity. To mitigate such distortion, we advocate for interpolation-free modeling, where models are designed to operate directly on asynchronously sampled and non-uniformly aligned signals, without relying on synthesized values that risk obscuring the true structure of the original signals.
(a) FFT Magnitude Spectrum and Distortion
(b) FFT Phase Spectrum and Delay
Figure 6: Spectral distortion caused by linear interpolation in the ETTh1 dataset. The left panel shows amplitude attenuation and spectral leakage in the magnitude spectrum, while the right panel illustrates phase delay in the FFT phase spectrum.
Appendix C Ablation Study
C.1 Additional Ablations
To complement the main experiments, we provide additional ablation studies highlighting the contribution of key architectural components in ChannelTokenFormer (CTF). As shown in Table 8, incorporating learnable channel-specific embeddings consistently improves performance across all datasets. This suggests that explicitly encoding channel identity through channel embeddings provides a strong inductive bias, particularly in heterogeneous multivariate settings.
Table 8: Effect of channel embedding on forecasting performance. Results are reported in CMSE, averaged over all prediction lengths. Overall, channel embedding improves performance on most datasets, indicating its effectiveness in modeling channel-specific dynamics.
Setting / Datasets ETT1 SolarWind Weather CHS
CTF with channel embedding 0.398 0.400 0.269 0.282
CTF w/o channel embedding 0.404 0.399 0.283 0.285
To further examine the effectiveness of our design beyond what was demonstrated in Section 5.2, we conduct an extended ablation study focusing on the three practical challenges of multivariate forecasting: Dependency, Asynchrony, and Missingness. As summarized in Table 9, we selectively disable each architectural component corresponding to the three core challenges addressed by our design, and evaluate all ablation settings under varying levels of test-time missingness. To remove Channel Dependence, we eliminate attention between channel tokens from different channels. To assess the impact of Dynamic Patching, we prevent the model from incorporating dominant frequency information extracted from FFT as well as channel-specific sampling periods. For Patch Masking, we disable the masking mechanism for block-wise missing intervals during both training and testing. Across all settings, the full model consistently achieves the best results, confirming that each architectural element contributes complementarily to the model’s overall robustness. Among these, patch masking emerges as critical for handling contiguous missing blocks, while channel dependency and dynamic patching offer additional gains under asynchronous and heterogeneous conditions.
Table 9: Results of the ablation study under channel-wise asynchronous sampling with block-wise test-time missing. Results are shown for different missing ratios $m$, with metrics averaged across all prediction lengths on the SolarWind dataset.
Test-time Missing Ratio $m=$ 0.125 $m=$ 0.250 $m=$ 0.375 $m=$ 0.500
Ablation Setting CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE
[…]
w/o Channel Dependence 0.412 0.464 0.435 0.484 0.457 0.507 0.482 0.532
w/o Dynamic patching 0.427 0.484 0.430 0.492 0.438 0.511 0.485 0.548
w/o Patch masking 0.435 0.486 0.486 0.525 0.536 0.564 0.574 0.593
C.2 Comparison of Attention Masking Strategies
We conduct an ablation study comparing six variants of attention masking strategies within our unified attention layer. To avoid analytical redundancy, all strategies share a common constraint: channel tokens are prohibited from attending to any other tokens within the same channel, including themselves. This restriction eliminates redundant intra-channel interactions and encourages richer cross-channel communication.
The six masking strategies fall into two categories: Channel-Independent (CI) and Channel-Dependent (CD). The key distinction lies in whether cross-channel attention among channel tokens is permitted. CD strategies are further subdivided based on the scope and indexing of allowed inter-channel connections.
  • Channel-Independent Masking (2 variants): All attention operations are confined to within each channel. No cross-channel interaction is allowed.
    • CI-ReadOnly: Channel tokens can attend to local tokens (read-only summarization), but not vice versa.
    • CI-Mutual: Local and channel tokens within a channel can attend to each other bidirectionally.
  • Channel-Dependent Masking (4 variants): Local tokens remain strictly within-channel, attending only to tokens from the same channel. In contrast, channel tokens are permitted to interact across channels, subject to specific masking rules.
    • Global Cross-Channel Token Attention: Each channel token can attend to all channel tokens from other channels, but not to those from its own channel. This facilitates rich cross-channel communication while maintaining strict inter-channel exclusivity.
        CD-ReadOnly: Extends CI-ReadOnly by allowing channel tokens to also attend to channel tokens from other channels. This allows each channel token to summarize both intra- and inter-channel information without influencing local token representations directly.
        CD-Mutual: Extends CI-Mutual with additional global attention among channel tokens across different channels. It enables two-way communication between channel tokens, potentially enhancing global coordination and information exchange across channels.
    • Indexed Cross-Channel Token Attention: When multiple channel tokens are assigned per channel, attention is restricted to those with matching indices across channels (e.g., index-0 tokens attend only to other index-0 tokens). This design encourages index-wise specialization and alignment.
        CD-ReadOnly-Indexed: Read-only variant allowing same-index channel tokens to attend to each other across channels. This selective communication enables controlled abstraction across semantically aligned channel positions.
        CD-Mutual-Indexed: Mutual variant enabling bidirectional intra-channel and same-index inter-channel attention. It balances local coherence and cross-channel consistency by unifying both local and indexed global interactions.
Figure 7: Visualization of the attention masks used in Channel-Independent (CI) masking strategies. CI-ReadOnly (left) allows only unidirectional attention from channel tokens to local tokens within the same channel, while CI-Mutual (right) permits bidirectional attention between local and channel tokens. In both cases, no cross-channel attention is allowed.
Figure 8: Visualization of the attention masks used in Channel-Dependent (CD) masking strategies with global cross-channel token attention. CD-ReadOnly (left) extends CI-ReadOnly by allowing channel tokens to attend to channel tokens from other channels, while CD-Mutual (right) permits bidirectional attention within channels and additionally enables global cross-channel attention among channel tokens.
Figure 9: Visualization of the attention masks used in Indexed Channel-Dependent (CD) masking strategies. CD-ReadOnly-Indexed (left) and CD-Mutual-Indexed (right) restrict cross-channel attention to channel tokens with the same index across different channels. This design promotes role-specific specialization while maintaining intra-channel locality for local tokens.
Table 10: Comparison of attention masking strategies on the SolarWind dataset under the block-wise test-time missing scenario. Results are reported in CMSE and CMAE, averaged over all prediction lengths with a missing ratio $m=0.125$. Among the methods, CD-ReadOnly and CD-ReadOnly-Indexed achieve the best performance under our proposed practical setting.
Strategies CI-ReadOnly CI-Mutual CD-ReadOnly CD-Mutual CD-ReadOnly-Indexed CD-Mutual-Indexed
CMSE 0.412 0.416 0.396 0.398 0.396 0.399
CMAE 0.464 0.468 0.458 0.458 0.458 0.460
Appendix D Forecasting Results in a Conventional Multivariate Setting
To demonstrate that our model remains effective even under standard multivariate scenarios, we additionally evaluate it on regularly sampled multivariate time series without any missing values. By aligning the sampling period and patch length across all channels, our model can be directly applied to this conventional setting. This evaluation shows that our model remains applicable beyond the proposed practical settings. For comparison, baseline results [4, 42, 39, 29, 15, 23, 34, 53, 7, 44, 50] are obtained either from the original publications or by reproducing them using the TSLib codebase [41]. We compare forecasting performance on five widely used benchmark datasets [55, 43]: ETTh1, ETTh2, ETTm1, ETTm2, and Weather. As shown in Table 11, our model achieves performance comparable to or better than state-of-the-art baselines across all benchmark datasets in the conventional forecasting setting.
Table 11: Full results of the long-term forecasting task in the conventional setting. The best results are highlighted in bold, and the second-best results are underlined.
Models
CTF(ours)
SimpleTM
TimeXer
TimeMixer
iTrans.
CrossGNN
RLinear
PatchTST
Crossformer
TiDE
TimesNet
DLinear
Metric
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
ETTh1
96
0.371
0.397
0.366
0.392
0.382
0.403
0.381
0.401
0.386
0.405
0.382
0.398
0.386
0.395
0.414
0.419
0.423
0.448
0.479
0.464
0.384
0.402
0.386
0.400
192
0.417
0.426
0.422
0.421
0.429
0.435
0.440
0.433
0.441
0.436
0.427
0.425
0.437
0.424
0.460
0.445
0.471
0.474
0.525
0.492
0.436
0.429
0.437
0.432
336
0.451
0.448
0.440
0.438
0.468
0.448
0.501
0.462
0.487
0.458
0.465
0.445
0.479
0.446
0.501
0.466
0.570
0.546
0.565
0.515
0.491
0.469
0.481
0.459
720
0.462
0.470
0.463
0.462
0.469
0.461
0.501
0.482
0.503
0.491
0.472
0.468
0.481
0.470
0.500
0.488
0.653
0.621
0.594
0.558
0.521
0.500
0.519
0.516
Avg
0.425
0.435
0.422
0.428
0.437
0.437
0.458
0.445
0.454
0.447
0.437
0.434
0.446
0.434
0.469
0.454
0.529
0.522
0.541
0.507
0.458
0.450
0.456
0.452
ETTh2
96
0.288
0.340
0.281
0.338
0.286
0.338
0.292
0.343
0.297
0.349
0.309
0.359
0.288
0.338
0.302
0.348
0.745
0.584
0.400
0.440
0.340
0.374
0.333
0.387
192
0.376
0.396
0.355
0.387
0.363
0.389
0.374
0.395
0.380
0.400
0.390
0.406
0.374
0.390
0.388
0.400
177
0.656
0.528
0.509
0.402
0.414
0.477
0.476
336
0.427
0.433
0.365
0.401
0.414
0.423
0.428
0.433
0.428
0.432
0.426
0.444
0.415
0.426
0.426
0.433
1.043
0.731
0.643
0.571
0.452
0.452
0.594
0.541
720
0.428
0.444
0.413
0.436
0.408
0.432
0.454
0.458
0.427
0.445
0.445
0.444
0.420
0.440
0.431
0.446
1.104
0.763
174
0.679
0.462
0.468
131
0.657
Avg
0.380
0.403
0.353
0.391
0.367
0.396
0.384
0.407
0.383
0.407
0.393
0.413
0.374
0.398
0.387
0.407
0.942
0.684
0.611
0.550
0.414
0.427
0.559
0.515
ETTm1
96
0.317
0.356
0.321
0.361
0.318
0.356
0.328
0.363
0.334
0.368
0.335
0.373
0.355
0.376
0.329
0.367
0.404
0.426
0.364
0.387
0.338
0.375
0.345
0.372
192
0.359
0.381
0.360
0.380
0.362
0.383
0.364
0.384
0.387
0.391
0.372
0.390
0.391
0.392
0.367
0.385
0.450
0.451
0.398
0.404
0.374
0.387
0.380
0.389
336
0.389
0.404
0.390
0.404
0.395
0.407
0.390
0.404
0.426
0.420
0.403
0.411
0.424
0.415
0.399
0.410
0.532
0.515
0.428
0.425
0.410
0.411
0.413
0.413
720
0.447
0.440
0.454
0.438
0.452
0.441
0.458
0.445
0.491
0.459
0.461
0.442
0.487
0.450
0.454
0.439
0.666
0.589
0.487
0.461
0.478
0.450
0.474
0.453
Avg
0.378
0.395
0.381
0.396
0.382
0.397
0.385
0.399
0.407
0.410
0.393
0.404
0.414
0.407
0.387
0.400
0.513
0.496
0.419
0.419
0.400
0.406
0.403
0.407
ETTm2
96
0.173
0.255
0.173
0.257
0.171
0.256
0.176
0.259
0.180
0.264
0.176
0.266
0.182
0.265
0.175
0.259
0.287
0.366
0.207
0.305
0.187
0.267
0.193
0.292
192
0.238
0.299
0.238
0.299
0.237
0.299
0.242
0.303
0.250
0.309
0.240
0.307
0.246
0.304
0.241
0.302
0.414
0.492
0.290
0.364
0.249
0.309
0.284
0.362
336
0.295
0.336
0.296
0.338
0.296
0.338
0.304
0.342
0.311
0.348
0.304
0.345
0.307
0.342
0.305
0.343
0.597
0.542
0.377
0.422
0.321
0.351
0.369
0.427
720
0.392
0.394
0.393
0.395
0.392
0.394
0.393
0.397
0.412
0.407
0.406
0.400
0.407
0.398
0.402
0.400
1.730
1.042
0.558
0.524
0.408
0.403
0.554
0.522
Avg
0.274
0.321
0.275
0.322
0.274
0.322
0.278
0.325
0.288
0.332
0.282
0.330
0.286
0.327
0.281
0.326
0.757
0.610
0.358
0.404
0.291
0.333
0.350
0.401
Weather
96
0.159
0.207
0.162
0.207
0.157
0.205
0.165
0.212
0.174
0.214
0.159
0.218
0.192
0.232
0.177
0.218
0.158
0.230
0.202
0.261
0.172
0.220
0.196
0.255
192
0.207
0.251
0.208
0.248
0.204
0.247
0.209
0.253
0.221
0.254
0.211
0.266
0.240
0.271
0.225
0.259
0.206
0.277
0.242
0.298
0.219
0.261
0.237
0.296
336
0.263
0.292
0.263
0.290
0.261
0.290
0.264
0.293
0.278
0.296
0.267
0.310
0.292
0.307
0.278
0.297
0.272
0.335
0.287
0.335
0.280
0.306
0.283
0.335
720
0.343
0.344
0.340
0.341
0.340
0.341
0.342
0.345
0.358
0.347
0.352
0.362
0.364
0.353
0.354
0.348
0.398
0.418
0.351
0.386
0.365
0.359
0.345
0.381
Avg
0.243
0.273
0.243
0.271
0.241
0.271
0.245
0.276
0.258
0.278
0.247
0.289
0.272
0.291
0.259
0.281
0.259
0.315
0.271
0.320
0.259
0.287
0.265
0.317
Appendix E Full Results of Our Proposed Practical Setting
Table 13 shows the results for Scenario 1 using linearly interpolated datasets, where practical data with channel-wise asynchronous sampling are converted to a regular sampling grid to accommodate baseline models. Table 14 presents the results for Scenario 2, which adds block-wise test-time missing inputs on top of asynchronous sampling. All results include error bars based on 5 random seeds to reflect performance variability. Across both scenarios, our model consistently demonstrates stronger performance than baseline methods.
Table 12: Full results of the Asynchronous long-term multivariate forecasting task (forward fill).
Strategy
Channel-Dependent
Channel-Independent
Models
CTF(ours)
TimeXer
TimeMixer
iTrans.
CrossGNN
TimesNet
PatchTST
DLinear
Metric
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
ETT1
192
0.336
0.371
0.358
0.386
0.348
0.377
0.355
0.386
0.363
0.382
0.371
0.397
0.344
0.376
0.354
0.378
336
0.363
0.390
0.392
0.412
0.379
0.395
0.394
0.411
0.398
0.402
0.418
0.423
0.376
0.398
0.388
0.397
768
0.427
0.430
0.440
0.440
0.449
0.435
0.463
0.450
0.460
0.439
0.464
0.458
0.447
0.439
0.448
0.436
1152
0.467
0.454
0.479
0.462
0.488
0.461
0.505
0.476
0.502
0.464
0.488
0.466
0.476
0.457
0.494
0.468
Avg
0.398
0.411
0.417
0.425
0.416
0.417
0.429
0.431
0.431
0.421
0.435
0.436
0.411
0.418
0.421
0.420
SolarWind
288
0.322
0.383
0.340
0.415
0.369
0.434
0.368
0.423
0.348
0.418
0.347
0.397
0.332
0.406
0.337
0.442
576
0.405
0.460
0.428
0.483
0.447
0.480
0.451
0.477
0.443
0.473
0.534
0.504
0.410
0.466
0.415
0.512
864
0.417
0.461
0.424
0.465
0.485
0.506
0.489
0.501
0.511
0.512
0.474
0.489
0.448
0.491
0.456
0.545
1152
0.456
0.488
0.489
0.514
0.491
0.522
0.544
0.526
0.572
0.547
0.553
0.525
0.461
0.490
0.469
0.557
Avg
0.400
0.448
0.420
0.469
0.448
0.486
0.463
0.482
0.468
0.488
0.477
0.479
0.413
0.463
0.419
0.514
Weather
144
0.179
0.225
0.222
0.257
0.204
0.243
0.235
0.271
0.196
0.240
0.253
0.285
0.200
0.237
0.207
0.261
288
0.242
0.275
0.261
0.285
0.277
0.296
0.288
0.307
0.265
0.311
0.261
0.288
0.292
0.320
0.260
0.308
576
0.311
0.322
0.331
0.331
0.348
0.346
0.349
0.343
0.326
0.339
0.335
0.332
0.351
0.355
0.320
0.356
864
0.343
0.347
0.361
0.357
0.352
0.355
0.380
0.371
0.366
0.368
0.377
0.366
0.372
0.359
0.355
0.395
Avg
0.269
0.293
0.294
0.308
0.295
0.310
0.313
0.323
0.289
0.314
0.307
0.318
0.304
0.318
0.286
0.330
CHS
120
0.104
0.062
0.129
0.070
0.111
0.062
0.125
0.074
0.106
0.075
0.137
0.074
0.135
0.066
0.111
0.097
240
0.165
0.087
0.177
0.089
0.177
0.089
0.184
0.098
0.168
0.099
0.190
0.099
0.216
0.094
0.186
0.171
480
0.312
0.140
0.321
0.142
0.318
0.139
0.329
0.144
0.310
0.159
0.340
0.148
0.335
0.142
0.375
0.288
840
0.548
0.213
0.559
0.214
0.572
0.216
0.570
0.216
0.549
0.224
0.602
0.224
0.614
0.230
0.695
0.435
Avg
0.282
0.126
0.296
0.129
0.294
0.127
0.302
0.133
0.283
0.139
0.317
0.136
0.325
0.133
0.342
0.248
Table 13: Full error bars for the channel-wise asynchronous long-term multivariate forecasting task.
Strategy Channel-Dependent Channel-Independent Models CTF(ours) TimeXer TimeMixer iTrans. CrossGNN TimesNet PatchTST DLinear Metric CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE ETT1 192 0.336 $\pm$ 0.002 0.371 $\pm$ 0.001 0.353 $\pm$ 0.001 0.381 $\pm$ 0.001 0.345 $\pm$ 0.001 0.372 $\pm$ 0.003 0.354 $\pm$ 0.003 0.383 $\pm$ 0.002 0.356 $\pm$ 0.001 0.373 $\pm$ 0.001 0.373 $\pm$ 0.010 0.397 $\pm$ 0.009 0.342 $\pm$ 0.001 0.373 $\pm$ 0.001 0.352 $\pm$ 0.001 0.373 $\pm$ 0.001 336 0.363 $\pm$ 0.001 0.390 $\pm$ 0.000 0.394 $\pm$ 0.001 0.410 $\pm$ 0.002 0.380 $\pm$ 0.002 0.394 $\pm$ 0.002 0.397 $\pm$ 0.001 0.411 $\pm$ 0.002 0.395 $\pm$ 0.001 0.397 $\pm$ 0.001 0.429 $\pm$ 0.011 0.428 $\pm$ 0.005 0.378 $\pm$ 0.002 0.397 $\pm$ 0.001 0.389 $\pm$ 0.001 0.396 $\pm$ 0.001 768 0.427 $\pm$ 0.002 0.430 $\pm$ 0.001 0.449 $\pm$ 0.003 0.443 $\pm$ 0.002 0.456 $\pm$ 0.002 0.437 $\pm$ 0.002 0.472 $\pm$ 0.003 0.454 $\pm$ 0.001 0.463 $\pm$ 0.002 0.438 $\pm$ 0.002 0.533 $\pm$ 0.026 0.483 $\pm$ 0.008 0.451 $\pm$ 0.004 0.440 $\pm$ 0.003 0.454 $\pm$ 0.003 0.438 $\pm$ 0.003 1152 0.467 $\pm$ 0.001 0.454 $\pm$ 0.003 0.484 $\pm$ 0.005 0.462 $\pm$ 0.001 0.498 $\pm$ 0.003 0.464 $\pm$ 0.004 0.514 $\pm$ 0.002 0.479 $\pm$ 0.003 0.507 $\pm$ 0.002 0.464 $\pm$ 0.003 0.535 $\pm$ 0.020 0.488 $\pm$ 0.012 0.485 $\pm$ 0.002 0.460 $\pm$ 0.002 0.502 $\pm$ 0.002 0.471 $\pm$ 0.003 SolarWind 288 0.322 $\pm$ 0.014 0.383 $\pm$ 0.020 0.337 $\pm$ 0.013 0.410 $\pm$ 0.012 0.369 $\pm$ 0.013 0.433 $\pm$ 0.021 0.367 $\pm$ 0.008 0.422 $\pm$ 0.005 0.347 $\pm$ 0.010 0.416 $\pm$ 0.013 0.347 $\pm$ 0.009 0.395 $\pm$ 0.006 0.326 $\pm$ 0.008 0.400 $\pm$ 0.005 0.336 $\pm$ 0.005 0.441 $\pm$ 0.005 576 0.405 $\pm$ 0.007 0.460 $\pm$ 0.010 0.427 $\pm$ 0.008 0.479 $\pm$ 0.010 0.447 $\pm$ 0.013 0.479 $\pm$ 0.013 0.450 $\pm$ 0.005 0.476 $\pm$ 0.004 0.443 $\pm$ 0.010 0.473 $\pm$ 0.010 0.534 $\pm$ 0.038 0.504 $\pm$ 0.026 0.410 $\pm$ 0.011 0.466 $\pm$ 0.006 0.415 $\pm$ 0.003 0.512 $\pm$ 0.003 864 0.417 $\pm$ 0.014 0.461 $\pm$ 0.014 0.426 $\pm$ 0.033 0.467 $\pm$ 0.027 0.489 $\pm$ 0.013 0.507 $\pm$ 0.014 0.488 $\pm$ 0.027 0.499 $\pm$ 0.014 0.511 $\pm$ 0.010 0.512 $\pm$ 0.007 0.474 $\pm$ 0.020 0.489 $\pm$ 0.013 0.447 $\pm$ 0.019 0.491 $\pm$ 0.017 0.457 $\pm$ 0.005 0.545 $\pm$ 0.004 1152 0.456 $\pm$ 0.010 0.488 $\pm$ 0.009 0.489 $\pm$ 0.017 0.514 $\pm$ 0.015 0.492 $\pm$ 0.023 0.523 $\pm$ 0.011 0.547 $\pm$ 0.011 0.527 $\pm$ 0.007 0.573 $\pm$ 0.012 0.547 $\pm$ 0.010 0.553 $\pm$ 0.010 0.526 $\pm$ 0.010 0.464 $\pm$ 0.009 0.494 $\pm$ 0.005 0.470 $\pm$ 0.001 0.557 $\pm$ 0.001 Weather 144 0.179 $\pm$ 0.003 0.225 $\pm$ 0.003 0.205 $\pm$ 0.007 0.245 $\pm$ 0.005 0.205 $\pm$ 0.017 0.240 $\pm$ 0.014 0.232 $\pm$ 0.003 0.271 $\pm$ 0.005 0.200 $\pm$ 0.007 0.248 $\pm$ 0.011 0.258 $\pm$ 0.022 0.288 $\pm$ 0.018 0.198 $\pm$ 0.014 0.236 $\pm$ 0.008 0.207 $\pm$ 0.001 0.261 $\pm$ 0.001 288 0.242 $\pm$ 0.003 0.275 $\pm$ 0.003 0.276 $\pm$ 0.004 0.291 $\pm$ 0.003 0.285 $\pm$ 0.017 0.300 $\pm$ 0.011 0.286 $\pm$ 0.002 0.305 $\pm$ 0.003 0.257 $\pm$ 0.005 0.296 $\pm$ 0.008 0.259 $\pm$ 0.021 0.285 $\pm$ 0.011 0.292 $\pm$ 0.023 0.320 $\pm$ 0.013 0.260 $\pm$ 0.001 0.308 $\pm$ 0.001 576 0.311 $\pm$ 0.002 0.322 $\pm$ 0.001 0.338 $\pm$ 0.003 0.335 $\pm$ 0.002 0.336 $\pm$ 0.016 0.332 $\pm$ 0.013 0.351 $\pm$ 0.001 0.346 $\pm$ 0.001 0.320 $\pm$ 0.006 0.336 $\pm$ 0.007 0.336 $\pm$ 0.007 0.335 $\pm$ 0.007 0.351 $\pm$ 0.016 0.355 $\pm$ 0.008 0.321 $\pm$ 0.001 0.356 $\pm$ 0.002 864 0.343 $\pm$ 0.003 0.347 $\pm$ 0.001 0.367 $\pm$ 0.004 0.360 $\pm$ 0.002 0.354 $\pm$ 0.004 0.353 $\pm$ 0.004 0.380 $\pm$ 0.002 0.368 $\pm$ 0.002 0.361 $\pm$ 0.009 0.372 $\pm$ 0.015 0.392 $\pm$ 0.023 0.380 $\pm$ 0.016 0.367 $\pm$ 0.003 0.359 $\pm$ 0.012 0.356 $\pm$ 0.005 0.395 $\pm$ 0.008 CHS 120 0.104 $\pm$ 0.002 0.062 $\pm$ 0.000 0.125 $\pm$ 0.005 0.067 $\pm$ 0.001 0.105 $\pm$ 0.004 0.061 $\pm$ 0.001 0.124 $\pm$ 0.007 0.072 $\pm$ 0.001 0.105 $\pm$ 0.005 0.070 $\pm$ 0.005 0.155 $\pm$ 0.009 0.079 $\pm$ 0.001 0.127 $\pm$ 0.007 0.066 $\pm$ 0.001 0.114 $\pm$ 0.003 0.106 $\pm$ 0.012 240 0.165 $\pm$ 0.002 0.087 $\pm$ 0.001 0.183 $\pm$ 0.006 0.091 $\pm$ 0.002 0.188 $\pm$ 0.016 0.090 $\pm$ 0.003 0.174 $\pm$ 0.007 0.092 $\pm$ 0.004 0.168 $\pm$ 0.000 0.102 $\pm$ 0.007 0.217 $\pm$ 0.010 0.104 $\pm$ 0.003 0.191 $\pm$ 0.009 0.092 $\pm$ 0.002 0.177 $\pm$ 0.004 0.132 $\pm$ 0.009 480 0.312 $\pm$ 0.004 0.140 $\pm$ 0.001 0.324 $\pm$ 0.004 0.139 $\pm$ 0.000 0.323 $\pm$ 0.029 0.141 $\pm$ 0.002 0.349 $\pm$ 0.008 0.149 $\pm$ 0.003 0.313 $\pm$ 0.001 0.151 $\pm$ 0.001 0.351 $\pm$ 0.016 0.154 $\pm$ 0.002 0.358 $\pm$ 0.008 0.145 $\pm$ 0.003 0.366 $\pm$ 0.027 0.269 $\pm$ 0.035 840 0.548 $\pm$ 0.008 0.213 $\pm$ 0.004 0.556 $\pm$ 0.019 0.216 $\pm$ 0.002 0.582 $\pm$ 0.016 0.216 $\pm$ 0.002 0.585 $\pm$ 0.009 0.217 $\pm$ 0.001 0.560 $\pm$ 0.004 0.239 $\pm$ 0.009 0.612 $\pm$ 0.035 0.229 $\pm$ 0.005 0.586 $\pm$ 0.006 0.226 $\pm$ 0.003 0.693 $\pm$ 0.068 0.435 $\pm$ 0.049
Table 14: Full error bars for the SolarWind dataset under block-wise test-time missingness with varying missing ratios $m$.
Strategy Channel-Dependent Channel-Independent Models CTF(ours) TimeXer TimeMixer iTrans. CrossGNN TimesNet PatchTST DLinear Metric CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE $m=0.125$ 288 0.331 $\pm$ 0.004 0.412 $\pm$ 0.003 0.355 $\pm$ 0.058 0.438 $\pm$ 0.050 0.372 $\pm$ 0.012 0.456 $\pm$ 0.023 0.368 $\pm$ 0.005 0.432 $\pm$ 0.005 0.358 $\pm$ 0.013 0.441 $\pm$ 0.012 0.353 $\pm$ 0.009 0.416 $\pm$ 0.006 0.372 $\pm$ 0.017 0.457 $\pm$ 0.010 0.352 $\pm$ 0.004 0.464 $\pm$ 0.004 576 0.393 $\pm$ 0.002 0.456 $\pm$ 0.001 0.434 $\pm$ 0.042 0.498 $\pm$ 0.038 0.452 $\pm$ 0.013 0.508 $\pm$ 0.013 0.451 $\pm$ 0.003 0.487 $\pm$ 0.003 0.445 $\pm$ 0.010 0.496 $\pm$ 0.008 0.494 $\pm$ 0.038 0.497 $\pm$ 0.026 0.454 $\pm$ 0.017 0.522 $\pm$ 0.012 0.428 $\pm$ 0.002 0.532 $\pm$ 0.003 864 0.418 $\pm$ 0.008 0.475 $\pm$ 0.007 0.444 $\pm$ 0.061 0.495 $\pm$ 0.017 0.474 $\pm$ 0.016 0.514 $\pm$ 0.022 0.481 $\pm$ 0.015 0.506 $\pm$ 0.006 0.505 $\pm$ 0.009 0.533 $\pm$ 0.005 0.474 $\pm$ 0.026 0.496 $\pm$ 0.016 0.486 $\pm$ 0.022 0.544 $\pm$ 0.016 0.466 $\pm$ 0.004 0.563 $\pm$ 0.004 1152 0.441 $\pm$ 0.005 0.489 $\pm$ 0.005 0.495 $\pm$ 0.032 0.533 $\pm$ 0.027 0.480 $\pm$ 0.027 0.519 $\pm$ 0.022 0.538 $\pm$ 0.006 0.534 $\pm$ 0.003 0.560 $\pm$ 0.011 0.566 $\pm$ 0.014 0.537 $\pm$ 0.011 0.529 $\pm$ 0.011 0.499 $\pm$ 0.019 0.538 $\pm$ 0.015 0.480 $\pm$ 0.001 0.575 $\pm$ 0.001 $m=0.25$ 288 0.347 $\pm$ 0.005 0.433 $\pm$ 0.003 0.386 $\pm$ 0.059 0.465 $\pm$ 0.048 0.389 $\pm$ 0.012 0.482 $\pm$ 0.023 0.387 $\pm$ 0.004 0.455 $\pm$ 0.004 0.375 $\pm$ 0.014 0.468 $\pm$ 0.013 0.371 $\pm$ 0.009 0.445 $\pm$ 0.006 0.430 $\pm$ 0.022 0.514 $\pm$ 0.011 0.374 $\pm$ 0.003 0.489 $\pm$ 0.004 576 0.404 $\pm$ 0.003 0.473 $\pm$ 0.001 0.469 $\pm$ 0.044 0.526 $\pm$ 0.036 0.472 $\pm$ 0.014 0.536 $\pm$ 0.013 0.472 $\pm$ 0.005 0.512 $\pm$ 0.004 0.455 $\pm$ 0.010 0.523 $\pm$ 0.008 0.523 $\pm$ 0.041 0.528 $\pm$ 0.025 0.533 $\pm$ 0.025 0.593 $\pm$ 0.018 0.447 $\pm$ 0.002 0.555 $\pm$ 0.002 864 0.434 $\pm$ 0.006 0.495 $\pm$ 0.005 0.496 $\pm$ 0.043 0.530 $\pm$ 0.012 0.484 $\pm$ 0.020 0.534 $\pm$ 0.026 0.504 $\pm$ 0.013 0.528 $\pm$ 0.008 0.509 $\pm$ 0.009 0.559 $\pm$ 0.004 0.481 $\pm$ 0.031 0.517 $\pm$ 0.017 0.573 $\pm$ 0.037 0.625 $\pm$ 0.021 0.482 $\pm$ 0.003 0.583 $\pm$ 0.003 1152 0.455 $\pm$ 0.008 0.508 $\pm$ 0.005 0.542 $\pm$ 0.021 0.564 $\pm$ 0.018 0.510 $\pm$ 0.023 0.536 $\pm$ 0.029 0.555 $\pm$ 0.006 0.557 $\pm$ 0.003 0.557 $\pm$ 0.013 0.590 $\pm$ 0.016 0.554 $\pm$ 0.016 0.554 $\pm$ 0.014 0.571 $\pm$ 0.043 0.607 $\pm$ 0.029 0.496 $\pm$ 0.001 0.596 $\pm$ 0.001 $m=0.375$ 288 0.363 $\pm$ 0.004 0.458 $\pm$ 0.002 0.412 $\pm$ 0.058 0.487 $\pm$ 0.048 0.421 $\pm$ 0.013 0.510 $\pm$ 0.023 0.415 $\pm$ 0.004 0.483 $\pm$ 0.005 0.400 $\pm$ 0.016 0.498 $\pm$ 0.014 0.398 $\pm$ 0.009 0.482 $\pm$ 0.008 0.492 $\pm$ 0.022 0.551 $\pm$ 0.016 0.400 $\pm$ 0.003 0.515 $\pm$ 0.003 576 0.418 $\pm$ 0.004 0.494 $\pm$ 0.003 0.498 $\pm$ 0.036 0.548 $\pm$ 0.025 0.509 $\pm$ 0.020 0.566 $\pm$ 0.015 0.501 $\pm$ 0.007 0.539 $\pm$ 0.004 0.473 $\pm$ 0.009 0.552 $\pm$ 0.007 0.538 $\pm$ 0.036 0.557 $\pm$ 0.020 0.612 $\pm$ 0.034 0.650 $\pm$ 0.020 0.470 $\pm$ 0.002 0.579 $\pm$ 0.002 864 0.445 $\pm$ 0.011 0.516 $\pm$ 0.006 0.542 $\pm$ 0.021 0.553 $\pm$ 0.008 0.511 $\pm$ 0.024 0.560 $\pm$ 0.026 0.536 $\pm$ 0.010 0.556 $\pm$ 0.008 0.521 $\pm$ 0.003 0.585 $\pm$ 0.004 0.499 $\pm$ 0.031 0.547 $\pm$ 0.016 0.651 $\pm$ 0.046 0.685 $\pm$ 0.030 0.501 $\pm$ 0.003 0.605 $\pm$ 0.002 1152 0.472 $\pm$ 0.008 0.532 $\pm$ 0.004 0.579 $\pm$ 0.013 0.587 $\pm$ 0.016 0.556 $\pm$ 0.023 0.565 $\pm$ 0.030 0.574 $\pm$ 0.009 0.584 $\pm$ 0.006 0.564 $\pm$ 0.021 0.615 $\pm$ 0.022 0.575 $\pm$ 0.028 0.585 $\pm$ 0.020 0.644 $\pm$ 0.061 0.667 $\pm$ 0.042 0.516 $\pm$ 0.001 0.618 $\pm$ 0.001 $m=0.5$ 288 0.390 $\pm$ 0.004 0.488 $\pm$ 0.003 0.445 $\pm$ 0.052 0.508 $\pm$ 0.056 0.456 $\pm$ 0.015 0.533 $\pm$ 0.022 0.460 $\pm$ 0.007 0.508 $\pm$ 0.005 0.429 $\pm$ 0.014 0.526 $\pm$ 0.011 0.421 $\pm$ 0.010 0.514 $\pm$ 0.008 0.566 $\pm$ 0.026 0.571 $\pm$ 0.012 0.428 $\pm$ 0.002 0.540 $\pm$ 0.003 576 0.444 $\pm$ 0.003 0.525 $\pm$ 0.002 0.528 $\pm$ 0.028 0.568 $\pm$ 0.015 0.542 $\pm$ 0.025 0.585 $\pm$ 0.014 0.550 $\pm$ 0.011 0.568 $\pm$ 0.004 0.496 $\pm$ 0.007 0.581 $\pm$ 0.005 0.574 $\pm$ 0.036 0.592 $\pm$ 0.018 0.674 $\pm$ 0.034 0.685 $\pm$ 0.022 0.494 $\pm$ 0.001 0.602 $\pm$ 0.002 864 0.470 $\pm$ 0.010 0.544 $\pm$ 0.005 0.586 $\pm$ 0.006 0.574 $\pm$ 0.013 0.548 $\pm$ 0.023 0.587 $\pm$ 0.023 0.597 $\pm$ 0.011 0.589 $\pm$ 0.006 0.538 $\pm$ 0.006 0.613 $\pm$ 0.007 0.528 $\pm$ 0.029 0.580 $\pm$ 0.014 0.708 $\pm$ 0.036 0.713 $\pm$ 0.023 0.522 $\pm$ 0.002 0.626 $\pm$ 0.001 1152 0.506 $\pm$ 0.006 0.563 $\pm$ 0.002 0.612 $\pm$ 0.018 0.607 $\pm$ 0.016 0.612 $\pm$ 0.024 0.594 $\pm$ 0.028 0.617 $\pm$ 0.011 0.615 $\pm$ 0.005 0.577 $\pm$ 0.035 0.642 $\pm$ 0.028 0.611 $\pm$ 0.038 0.621 $\pm$ 0.023 0.704 $\pm$ 0.071 0.703 $\pm$ 0.037 0.538 $\pm$ 0.001 0.640 $\pm$ 0.001
Table 15: Average forecasting performance of our CTF model on the SolarWind dataset, evaluated under block-wise missing at test time with varying missing ratios $m$, and trained with different patch masking ratios $t$.
Train missing ratio
$t$ = 0.1
$t$ = 0.2
$t$ = 0.3
$t$ = 0.4
$t$ = 0.5
Metric
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
$m=0.125$
0.403
0.462
0.397
0.459
0.400
0.461
0.396
0.458
0.401
0.462
$m=0.250$
0.451
0.500
0.427
0.485
0.427
0.485
0.410
0.477
0.412
0.478
$m=0.375$
0.461
0.525
0.438
0.508
0.434
0.505
0.424
0.500
0.428
0.499
$m=0.500$
0.543
0.581
0.488
0.548
0.476
0.540
0.452
0.530
0.456
0.528
Table 16: Full results of the ETT1 dataset under short-range test-time missing intervals with varying missing ratios $m$. CTF is trained with a random patch masking ratio of 0.4 to enhance robustness against missing inputs.
Strategy
Channel-Dependent
Channel-Independent
Models
CTF(ours)
TimeXer
TimeMixer
iTrans.
CrossGNN
TimesNet
PatchTST
DLinear
Metric
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
$m=0.1$
192
0.353
0.385
0.373
0.397
0.368
0.391
0.380
0.405
0.379
0.393
0.399
0.420
0.369
0.396
0.382
0.398
336
0.383
0.405
0.412
0.425
0.401
0.412
0.419
0.429
0.416
0.415
0.454
0.448
0.405
0.419
0.419
0.420
768
0.445
0.442
0.469
0.460
0.474
0.453
0.489
0.468
0.478
0.452
0.550
0.501
0.475
0.460
0.481
0.460
1152
0.480
0.463
0.505
0.480
0.515
0.479
0.531
0.493
0.520
0.476
0.556
0.505
0.507
0.478
0.528
0.491
Avg
0.415
0.424
0.440
0.441
0.440
0.434
0.455
0.449
0.448
0.434
0.490
0.468
0.439
0.438
0.453
0.442
$m=0.2$
192
0.373
0.400
0.397
0.417
0.395
0.412
0.407
0.427
0.405
0.414
0.423
0.441
0.397
0.419
0.415
0.423
336
0.402
0.419
0.434
0.443
0.426
0.431
0.442
0.448
0.441
0.435
0.472
0.466
0.433
0.441
0.450
0.444
768
0.463
0.456
0.496
0.481
0.497
0.471
0.508
0.485
0.499
0.469
0.565
0.516
0.498
0.477
0.511
0.482
1152
0.498
0.478
0.535
0.503
0.535
0.496
0.550
0.510
0.537
0.492
0.571
0.519
0.532
0.496
0.556
0.512
Avg
0.434
0.438
0.466
0.461
0.463
0.453
0.477
0.467
0.471
0.452
0.508
0.486
0.465
0.458
0.483
0.465
$m=0.3$
192
0.396
0.413
0.426
0.436
0.426
0.434
0.436
0.446
0.435
0.434
0.447
0.460
0.427
0.441
0.449
0.445
336
0.422
0.431
0.459
0.460
0.454
0.449
0.469
0.466
0.468
0.454
0.491
0.482
0.462
0.462
0.484
0.466
768
0.481
0.469
0.527
0.503
0.521
0.489
0.527
0.498
0.521
0.485
0.584
0.532
0.520
0.493
0.540
0.501
1152
0.515
0.491
0.567
0.525
0.556
0.513
0.567
0.522
0.557
0.507
0.588
0.533
0.556
0.514
0.584
0.531
Avg
0.453
0.451
0.495
0.481
0.489
0.471
0.500
0.483
0.495
0.470
0.527
0.502
0.491
0.478
0.514
0.486
$m=0.4$
192
0.414
0.426
0.451
0.454
0.455
0.454
0.462
0.463
0.462
0.453
0.474
0.480
0.456
0.463
0.480
0.467
336
0.440
0.443
0.486
0.479
0.481
0.468
0.492
0.481
0.495
0.472
0.511
0.497
0.491
0.483
0.514
0.487
768
0.499
0.481
0.562
0.527
0.548
0.508
0.548
0.512
0.545
0.502
0.608
0.549
0.542
0.510
0.570
0.521
1152
0.533
0.503
0.602
0.547
0.579
0.530
0.589
0.537
0.579
0.523
0.609
0.549
0.580
0.531
0.612
0.549
Avg
0.471
0.463
0.526
0.502
0.516
0.490
0.523
0.499
0.520
0.487
0.550
0.519
0.517
0.497
0.544
0.506
$m=0.5$
192
0.437
0.440
0.482
0.475
0.490
0.477
0.493
0.483
0.494
0.474
0.505
0.501
0.490
0.487
0.516
0.489
336
0.462
0.457
0.518
0.501
0.512
0.489
0.520
0.499
0.525
0.492
0.537
0.515
0.525
0.507
0.549
0.509
768
0.520
0.495
0.607
0.555
0.577
0.527
0.570
0.526
0.572
0.520
0.634
0.566
0.568
0.528
0.602
0.540
1152
0.552
0.517
0.647
0.573
0.604
0.547
0.610
0.550
0.604
0.540
0.629
0.563
0.607
0.549
0.642
0.567
Avg
0.493
0.477
0.563
0.526
0.546
0.510
0.548
0.515
0.549
0.506
0.576
0.536
0.547
0.518
0.577
0.526
Appendix F Forecast Visualization
Figure 10(a) illustrates the forecasts of four models including our CTF for channel 3 in the ETT1 dataset under Scenario 1 in Section 5. The prediction horizon shows periodic sharp drops and recoveries, which are critical for accurate forecasting. Our model closely tracks these rapid transitions, while baseline models tend to underpredict their magnitude. We attribute this to the interpolation-free nature of our model, which avoids the frequency biases introduced by linear interpolation.
We also examine Scenario 2 in Section 5, which involves both asynchronous sampling and block-wise test-time missing. Figure 10(b) presents a visualization of model predictions for the solar power channel from the SolarWind dataset. In this sample, a missing block of length 144 is present in the middle of the test input. This causes TimeMixer and TimeXer to produce noticeably noisy forecasts. Although CrossGNN shows relatively higher robustness, its predictions exhibit a drift in the overall trend and an underestimation of the periodic transitions. In contrast, our model accurately identifies the missing region and applies masking, effectively preventing the injection of unreliable inputs. As a result, it successfully captures both the timing and magnitude of the periodic transitions, even across the missing block.
(a) Channel 3 in the ETT1 dataset under asynchronous sampling. Our model accurately captures periodic sharp drops and recoveries, whereas the baselines tend to underpredict their magnitude.
(b) Solar power channel in the SolarWind dataset under asynchronous sampling with block-wise test-time missing. Despite a long missing block, our model robustly captures periodic transitions, while the baselines exhibit noise or trend drift.
Figure 10: Visualization of forecasting results on (a) the ETT1 and (b) the SolarWind datasets.
Appendix G Discussion
To jointly capture temporal dynamics and cross-channel dependencies under the practical settings, we proposed a unified attention masking strategy that merges intra- and inter-channel attention into a single attention layer. This design avoids the need for separate attention modules, reducing the number of parameters and simplifying the overall model architecture. This parameter efficiency is particularly beneficial for datasets with limited data, as it reduces the risk of overfitting while maintaining the model’s ability to capture essential temporal and cross-channel patterns. However, this unified structure introduces scalability challenges when applied to high-dimensional datasets. As the number of channels increases, the attention layer must process a quadratic number of token interactions within a single computation step, leading to substantial computational overhead and slower training. A promising direction to mitigate this issue is to optimize the attention computation, for example, by decomposing the attention map or distributing the computation across multiple smaller attention modules.
Beyond scalability, another limitation lies in the objective function. Although our model demonstrates strong predictive performance across diverse practical settings, it still suffers from amplitude attenuation—especially over longer forecasting horizons—due to the use of a standard mean squared error loss. To address this, it is crucial to explore alternative loss functions that better preserve spectral characteristics and periodicity, rather than relying solely on point-wise accuracy. This issue also extends to evaluation. Conventional metrics such as MSE may fail to penalize overly smoothed forecasts, often favoring models that suppress meaningful high-frequency or transient patterns. As a result, structurally inaccurate predictions may still appear favorable under MSE, particularly in long-term forecasting scenarios dominated by coarse trends. Refining evaluation metrics to better capture structural fidelity is therefore essential. Designing both loss functions and evaluation protocols that align with the characteristics of asynchronous, partially observed time series remains an important and open research challenge.